2. Given the following table:+ Net high low Stock Ticker Div % P/E 00s High Low Close chg 2.5 3831 42.00 39.51 4150-1.90 162 6.09 5.90 6.09 +0.12 52W 52W Yield Vol 45.39 19.75 ResMed RMD 1.63 3.55 Revion A REV 7 25 55.13 RioTinto RTP 2.30 3.2 31.31 16.63 RitchieBr RBA 844 1.75 RiteAld RAD 38.63 18.81 RobtHalfRH 168 72.75 7184 72.74 +0.0 20.9 15 24.49 24.29 24.49-0.01 31028 450420 431 +021 26.5 6517 27.15 26.50 26.50 +0.14 5125 27.69 Rockwell ROK 1.02 2.1 14.5 6412 47.99 47.00 4754 +0.24 a. How many share of Rockwell could you purchase with $25,000? b. What would be your annual dividend income from those shares?+ c. What is the firms current earnings per share (EPS)?+ d. What was the firms closing price yesterday?+
Answer #1

1.
Current Price=47.54
So, number of shares thatcan be purchased for 25000=25000/47.54=525.8729=525
2.
Annual dividends=525*1.02=535.5
3.
Current EPS=Price/(P/E)=47.54/14.5=3.278621
4.
CLosing price yesterday=closing price today-net change=47.54-0.24=47.30